OpenStudy (anonymous):

condense into single logarithmeic expression 4log(x)+log(z)-log (sqrtw)

OpenStudy (anonymous):

x^4*z/sqrt(w)

OpenStudy (anonymous):

\[x ^{4}z/\sqrt{w}\]

OpenStudy (anonymous):

\[ log\frac{x^4z}{\sqrt{w}}\]